OpenStudy (anonymous):

A population of horses has individuals that are either gray or chestnut colored. The gray color allele, G, is dominant over the chestnut color allele, g. If the frequency of the gray allele in the population is 0.8, what percent of horses are gray? A.) 64 percent B.) 96 percent C.) 32 percent D.) 80 percent

OpenStudy (anonymous):

1 = p^2 + 2pq + q^2 this is the equation you have to use

OpenStudy (matt101):

If the frequency of G = 0.8, the frequency of g = 0.2. In the equation you posted, p and q are general terms that can refer to any pair of alleles, so in this case we'll use G and g instead of p and q. Since gray, G, is the dominant allele, all horses with genotype GG (homozygous dominant) or Gg (heterozygous) will be gray, so you just have to add the two proportions. Percent homozygous dominant = 0.8^2 = 0.64 Percent heterozygous = 2*0.8*0.2 = 0.32 0.64 + 0.32 = 0.96 Alternatively, just subtract the proportion of horses that are NOT gray from 1, which will give you the proportion of horses that are: 1 - 0.2^2 = 1 - 0.04 = 0.96

OpenStudy (anonymous):

thanks nice job !!! follow me!